# references.bib

### v4

@@ 3.6 The latent states in a dynamic symbolic policy are transparent
Across the experiments, GP was able to find successful dynamic symbolic policies. The best validation fitness obtained with each method is presented in Table 4h. GP-D is generally competitive with black-box NDEs, and outperforms the NDE in 3 out of 7 experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. The dynamic symbolic policies compete with the black-box baseline, while consisting of interpretable mathematical expressions. The best static and dynamic symbolic policies are displayed in Table 1. Visualising the dynamic symbolic policies reveals that the latent states exhibits transparent properties that allow us to investigate how the memory is used to solve control tasks.
First, we investigated the functionality of the best policy evolved to stabilise the SHO with varying parameters settings (Table 1, SHO, Varying parameters, Dynamic). Our interpretation of the symbolic expression of variable $a_{1}$ is that it computes the distance between the target and position, adds the velocity and subtracts itself to decrease the influence of the noisy observations. Variable $a_{2}$ integrates $a_{1}$ multiplied with a constant, which we interpret as the estimate of the cumulative error. The readout subtracts $a_{2}$ and the cube of $a_{1}$ from the target state to output the policy’s control.
To demonstrate the generalisability of this policy within a single trial, we change the parameters of the SHO during the experiment (Fig. 5a, top row). The policy manages to keep the state close to the target state by increasing the strength of its negative control signal (Fig. 5a, middle row). Variable $a_{1}$ initially fluctuates until the position is stabilised close to the target, where after it stays close to the origin (Fig. 5a, bottom row, red curve). Variable $a_{2}$ first decreases, but increases during the trajectory to adjust to the changes in the environment, which causes the control force to increase (Fig. 5a, bottom row, blue curve).
The latent state should remain transparent when a policy becomes non-linear to still offer interpretability in more complex tasks than the SHO. To demonstrate the transparency in non-linear policies, the best policy obtained from acrobot swing-up task with observation noise is analysed (Table 1, Acrobot, Observation noise, Dynamic). Variable $a_{1}$ computes the difference between the angles of the two links. Variable $a_{2}$ inputs the angle of the first link in two nested sine functions, multiplied with a scalar. The readout subtracts the cosine of $a_{1}$ from $a_{2}$. Our interpretation of this policy is that it applies force in the direction of the first link compared to the centre point to build momentum, and increases the force when the two links are not closely aligned to add extra momentum to the second link, resembling a bang-bang controller.
The mathematical expressions of the non-linear policy are less straightforward to interpret than the linear policy. Analysing the latent state of this policy helps with understanding the functioning of this policy. The policy quickly accomplishes the swing up of the acrobot (Fig. 5b, top row). The control force applied by the policy shows expected behaviour, switching between -1 and 1 to build momentum, and remains constant longer when the second link is close to the threshold height (Fig. 5b, middle row). The latent state shows similar behaviour, although variable $a_{1}$ fluctuates with a higher amplitude (Fig. 5b, bottom row). The latent states are not as expressive as in the policy in Fig. 5a, and it is difficult to understand the individual contribution of each variable. Nonetheless, there is still more transparency in these dynamics than in a black-box policy.
@@
For application purposes, it is also relevant that policies remain interpretable in an industrial setting, such as the CSTR. The best policy discovered for stabilising the CSTR is analysed for transparency (Table 1 CSTR, Dynamic). We interpret the state equation of variable $a_{1}$ as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable $a_{2}$ computes the error between the temperature and setpoint, and adds a self-connection to improve robustness to noise. The error estimate is scaled by the sum of the control, $a_{2}$ and a constant. The readout scales $a_{2}$ with the squared of $a_{1}$, and adds a multiplication of $a_{1}$ with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, scaled increasingly during the trajectory.
To increase the number of responses in the trial, the setpoint temperature is changed several times during the trajectory. The policy efficiently stabilises the temperature at each setpoint temperature (Fig. 5c, top row). The control signal reacts when the setpoint temperature changes, and remains close to zero when the temperature has to increase. The concentration of the reactant in the main reactor goes to zero quickly and only slightly increases when the control input changes (Fig. 5c, middle row). Variable $a_{1}$ converges to 2.92 as expected (Fig. 5c, bottom row, red curve), while $a_{2}$ shows large responses when the setpoint temperature changes (Fig. 5c, bottom row, blue curve). Together, the latent variables allow the policy to change its control output accordingly.
@@
These example trials demonstrate that the dynamic symbolic policies are not only interpretable, but also offer transparency in the latent states. The latent state variables vary in their level of expressiveness, but overall analysing the latent state aids with understanding the functionality of a policy. The latent state showed to be beneficial for the robustness of a policy, however it also further improves the interpretability of symbolic policies.
@@ 4 Discussion
In this paper, we discovered interpretable and high-performing white-box control systems to solve control problems. The conventional approach is to find static symbolic policies, but these are difficult to extend to real-world challenges like partial observability and volatility. To improve the robustness of control systems, we evolved dynamic symbolic policies with memory with genetic programming (GP). The dynamic policies successfully solved control tasks with noisy and partial state observations, where the static policies failed to solve the task. Furthermore, dynamic symbolic policies generalise better to tasks with varying environment settings than static policies. These results show the benefits of the memory in the dynamic symbolic policies, and these policies can still be evolved efficiently with GP in terms of generations and population size. The symbolic policies were compared with a neural differential equation optimised with covariance-matrix adaptation evolution strategies as a proven black-box baseline [salimans2017evolution]. The results demonstrate that dynamic symbolic policies are evolved that generally obtain comparable performance as the black-box baseline, although competitive symbolic policies are not consistently found across evolutionary runs.
Besides adequately solving control tasks, the dynamic symbolic policies have other advantages, such as that they consist of interpretable expressions and display transparent dynamics, as demonstrated in Fig. 5. These features help understanding the functioning of control systems, therefore contributing to identification of undesired behaviour. Additionally, the interpretability allows us to learn new approaches to solve tasks from the policies. To further increase the interpretability in the dynamic symbolic policies, the readout could be learned as a numerical vector that maps the latent state to a control signal. This way, all non-linear functionality is evolved in the latent state, and the relation between the latent state and control signal becomes easier to understand, which is currently not always trivial (Table 1). The consistency of evolution also improves, as less trees have to be optimised with GP, although another method for optimising the readout vector should be introduced.
@@
Besides introducing the concept of dynamic symbolic policies, this research also resulted in a general GP framework that is not limited to optimisation of dynamic symbolic policy. First of all, other structures of symbolic policies can be configured, with the possibility to explicitly define different types of trees with unique inputs and functionality in the policy. Furthermore, the framework could also be used problems outside control, for example symbolic regression [bongard2007automated], discovering learning rules [jordan2021evolving] or other directions within scientific discovery [wang2023scientific, vinuesa2024opportunities, rackauckas2020universal].
The memory in dynamic policies demonstrated to improve the generalisability to different parameter settings in an environment. However, this could be further extended to evolve a single policy that can solve different control problems, which relates to meta-learning. In [kelly2021evolving], genetic programming is used to construct a memory-based agent that solves six different control problems under partial state observability, which has to recognize the task at hand purely from the observations. Our work could be integrated to replace the model of the agent with a dynamic symbolic policy. The interpretable policy expression and transparent latent state would aid in understanding how the agent solves different tasks with the same policy.
@@ Appendix A Genetic programming
Our implementation of genetic programming followed Algorithm 1. The hyperparameters and function sets we used in each experiment are presented in Table 2. The choice for the population size and number of generations (see Section 2.4) were chosen to balance efficiency and consistent convergence empirically. The elite percentage was set to 0.1 in every experiment. The choice of the function sets are explained in the corresponding environment sections.
@@
Table 2: Hyperparameters and function sets of the genetic programming algorithm used in every experiment.
@@
Appendix B Acrobot parameters
The acrobot dynamics are defined in terms of the variables
$d_{1}$ $=m_{1}l_{c1}^{2}+m_{2}\left(l_{1}^{2}+l_{c2}^{2}+2l_{1}l_{c2}\cos\left(\theta_{2}\right)\right)+I_{1}+I_{2}$
$d_{2}$ $=m_{2}\left(l_{c2}^{2}+l_{1}l_{c2}\cos\left(\theta_{2}\right)\right)+I_{2}$
$\phi_{1}$ $=-m_{2}l_{1}l_{c2}\dot{\theta}_{2}^{2}\cos\left(\theta_{2}\right)-2m_{2}l_{1}l_{c2}\dot{\theta}_{2}\dot{\theta}_{1}\sin\left(\theta_{2}\right)+\left(m_{1}l_{c1}+m_{2}l_{1}\right)g\cos\left(\theta_{1}-\frac{\pi}{2}\right)+\phi_{2}$
$\phi_{2}$ $=m_{2}l_{c2}g\cos\left(\theta_{1}+\theta_{2}-\frac{\pi}{2}\right)$
with parameters of the system as shown in Table 3.
@@ Appendix B Acrobot parameters
Appendix C Continuous stirred tank reactor parameters
The descriptions of each parameter in the CSTR dynamics are shown in Table 4. Table 4 also contains the value, initial condition or sample range for each parameter. The variables that are indicated with a range are uniformly sampled within this range in every trajectory. The initial conditions of $c$, $T_{r}$ and $T_{c}$ are uniformly sampled from the given distribution. The controlled parameter is clipped between 0 and 300. The reaction rate is defined as $k(T_{r})=k_{0}\exp\left(-{E}/{RT_{r}}\right)$.

### v5

@@ 4.6 The latent variables in dynamic symbolic policies are transparent
Across the experiments, GP was able to find successful dynamic symbolic policies. The best validation fitness obtained with each method is presented in Table 1. GP-D is generally competitive with black-box NDEs, and outperforms the NDE in 3 out of 7 experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. Table 1 also presents the complexity of the best policies in terms of the number of mathematical operations, variables and constants. GP-S always obtains policies with lower complexity than GP-D, which is partially accounted for by the fact that GP-D consists of three trees. The dynamic symbolic policies have an increased complexity compared to the static policies, but this complexity allows the policies to obtain higher performance in challenging experiments. Compared to the NDE, the complexity of GP-D is still significantly lower, while demonstrating competitive performance and more interpretability. The best static and dynamic symbolic policies are displayed in Table 2. Visualising the dynamic symbolic policies reveals that the latent variables exhibits transparent properties that allow us to investigate how the memory is used to solve control tasks.
@@
For this purpose, the best dynamic symbolic policy discovered with GP for stabilising the CSTR is analysed for its control and transparency (Table 2 CSTR, Dynamic). We interpret the state equation of variable $a_{1}$ as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable $a_{2}$ computes the error between the temperature and setpoint, and adds a self-connection to improve robustness to noise. The error estimate is scaled by the sum of the control, $a_{2}$ and a constant. The readout scales $a_{2}$ with the squared of $a_{1}$, and adds a multiplication of $a_{1}$ with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, scaled increasingly during the trajectory.
This policy is subjected to a set of challenging experiments on the CSTR with varying parameter settings that were not included during evolution to demonstrate its generalisability and the flexibility of the latent state. During evolution, the setpoint temperature was fixed within a trajectory, but in Fig. 5b, the setpoint changes throughout the trajectory to test whether the policy can adjust its control. At the start of the trajectory, the temperature in the main reactor is quickly stabilised around the setpoint temperature without requiring much control (Fig. 5b, top row). However, when the setpoint temperature is changed, the policy accordingly adjusts its control signal to decrease the temperature to the setpoint. Similarly, the policy can successfully stabilise the temperature at each setpoint temperature. Latent variable $a_{1}$ converges to 2.92 as expected (Fig. 5b, bottom row, red curve), but $a_{2}$ shows large responses when the setpoint temperature changes (Fig. 5c, bottom row, blue curve), which is indirectly reflected in the control signal.
In industrial settings, not only the target may vary, but the environment itself can be volatile. To demonstrate the robustness of the dynamic symbolic policy, the temperature of the reactant that flows into the main reactor increases throughout the trajectory, which requires the policy to continuously adapt to satisfy the objective (Fig. 5c). The temperature in the main reactor increases constantly, but the policy increases its control signal, which consequently decreases the temperature in the cooling jacket and keeps the temperature in the reactor close to the setpoint (Fig. 5c, top row). Latent variable $a_{2}$ also increases over time, showing it reacts to the changing environment (Fig. 5c, bottom row). When the setpoint temperature is altered halfway, the latent state adjusts, and the control is temporarily decreased, after which is steadily increases again to match the changing feed temperature.
The previous two experiments tested for robustness to changing targets or environment settings. Another realistic scenario is that sensors deteriorate, and the observations may become less accurate. In Fig. 5d, the noise applied to the observations of the temperature in the reactor and coolant is increased by a factor of 4 compared to the setting during evolution. The policy still manages to stabilise the temperature around the setpoint, even after a change of the setpoint temperature (Fig. 5d, top row). The latent variable $a_{2}$ behaves more noisy (Fig. 5d, bottom row), which is also reflected in the control signal.
Finally, an external sinusoidal influence is added to the dynamics of the temperature in the reactor (Fig. 5e). The sinusoid resembles the oscillating temperature outside of the reactor during day and night. Although the policy stabilises the temperature around the setpoint, the oscillations in the temperature are still visible (Fig. 5e, top row). The control signal and the temperature in the cooling jacket show larger oscillations, which means the policy has to heavily adjust to compensate for the external influence. The latent variable $a_{2}$ also exhibits oscillatory behaviour, with a different pattern occurring after the setpoint has been changed (Fig. 5e, bottom row).
@@
These experiments demonstrate that the dynamic symbolic policies can generalise to challenging environment settings that were not used during evolution. Additionally, the equations of the dynamic symbolic policy are interpretable, and the latent variables offer transparency in the use of memory. The latent variables vary in their level of expressiveness, but overall analysing the latent variables aids with understanding the functionality of a policy.
@@ 5 Discussion
In this paper, we discovered interpretable and high-performing white-box control systems to solve control problems. The conventional approach is to find static symbolic policies, but these are difficult to extend to real-world challenges like partial observability and volatility. To improve the robustness of control systems, the policies have to be extended with memory. Before, memory was added to symbolic policies by providing time-lagged observations or allowing the policy to explicitly store and retrieve data. Here, we extend the policy with latent variables that are governed by differential equations. The latent variables continuously integrate observations in their dynamics, which allows for a more scalable and general approach compared to using time-lagged observations. By evolving symbolic state equations of the latent variables with genetic programming (GP), the resulting policies consist of interpretable equations and the memory is transparent.
The dynamic policies successfully solved control tasks with noisy and partial state observations, where the static policies obtain limited results. Furthermore, dynamic symbolic policies generalise better to tasks with varying environment settings than static policies. These results show the benefits of the memory in the dynamic symbolic policies, and these policies can still be evolved efficiently with GP in terms of generations and population size. The symbolic policies were compared with a neural differential equation optimised with covariance-matrix adaptation evolution strategies as a proven black-box baseline [salimans2017evolution]. The results demonstrate that dynamic symbolic policies are evolved that generally obtain comparable performance as the black-box baseline, although competitive symbolic policies are not consistently found in every evolutionary run.
Besides adequately solving control tasks, the dynamic symbolic policies have other advantages, such as that they consist of interpretable expressions and display transparent dynamics, as demonstrated in Fig. 5. These features help understanding the functioning of control systems, therefore contributing to identification of undesired behaviour. Additionally, the interpretability allows us to learn new approaches to solve tasks from the policies. To further increase the interpretability in the dynamic symbolic policies, the readout could be learned as a linear numerical vector that maps the latent variables to a control signal. This way, all non-linear functionality is evolved in the latent state, and the relation between the latent variables and control signal becomes easier to understand, which is currently not always trivial (Table 2). The consistency of evolution also improves, as less trees have to be optimised with GP, although another method for optimising the readout vector should be introduced.
@@
Besides introducing the concept of symbolic policies with dynamic latent memory, this research also resulted in a general GP framework that is not limited to optimisation of dynamic symbolic policies. First of all, other structures of symbolic policies can be configured, with the possibility to explicitly define different types of trees with unique inputs and functionality in the policy. Furthermore, the framework could also be used problems outside control, for example symbolic regression [bongard2007automated], discovering learning rules [jordan2021evolving] or other directions within scientific discovery [wang2023scientific, vinuesa2024opportunities, rackauckas2020universal].
@@ Appendix A Genetic programming
Our implementation of genetic programming follows Algorithm 1. The hyperparameters and function sets we used in each experiment are presented in Table 3. The choice for the population size and number of generations (see Section II-D) were chosen to balance efficiency and consistent convergence empirically. The elite percentage was set to 0.1 in every experiment. The choice of the function sets are explained in the corresponding environment sections.
@@
Table 3: Hyperparameters and function sets of the genetic programming algorithm used in each experiment.
@@
Appendix B Stochastic Harmonic Oscillator
The dynamics of the stochastic harmonic oscillator (SHO) are given by
$d\mathbf{x}=(\mathbf{Ax}+\mathbf{b}{u})dt+\mathbf{v}d{w}$ (8)
with
with $\mathbf{Q}=\textrm{diag}(0.5,0)$ and $r=0.5$.
Here, $\mathbf{Q}$ punishes the policy for the distance between $\mathbf{x}$ and some target state $\mathbf{x}^{\ast}$ and $r$ penalises the control strength. The initial condition $\mathbf{x}_{0}$ is randomly sampled from a normal distribution with zero mean and variance of 3 and 1 for the position and velocity respectively. The target position is chosen randomly by sampling uniformly between -3 and 3 in each trial, while the target velocity is always set to zero.
Appendix C Acrobot
The dynamics of the acrobot are defined as in [sutton1995generalization]. Let $\theta_{1}$ and $\theta_{2}$ denote the angles of the first and second link, respectively. Their dynamics are given by
$\ddot{\theta}_{1}$ $=-\frac{d_{2}\ddot{\theta}_{2}+\phi_{1}}{d_{1}}$ (11)
$\ddot{\theta}_{2}$ $=\frac{u_{1}+d_{1}^{-1}d_{2}\phi_{1}-m_{2}l_{1}lc_{2}\dot{\theta}_{1}^{2}\sin(\theta_{2})-\phi_{2}}{m_{2}l_{c2}^{2}+I_{2}-d_{1}^{-1}d_{2}^{2}}$ (12)
with
$d_{1}$ $=m_{1}l_{c1}^{2}+m_{2}\left(l_{1}^{2}+l_{c2}^{2}+2l_{1}l_{c2}\cos\left(\theta_{2}\right)\right)+I_{1}+I_{2}$ (13)
$d_{2}$ $=m_{2}\left(l_{c2}^{2}+l_{1}l_{c2}\cos\left(\theta_{2}\right)\right)+I_{2}$ (14)
$\phi_{1}$ $=-m_{2}l_{1}l_{c2}\dot{\theta}_{2}^{2}\cos\left(\theta_{2}\right)-2m_{2}l_{1}l_{c2}\dot{\theta}_{2}\dot{\theta}_{1}\sin\left(\theta_{2}\right)$ (15)
$+\left(m_{1}l_{c1}+m_{2}l_{1}\right)g\cos\left(\theta_{1}-\frac{\pi}{2}\right)+\phi_{2}$ (16)
$\phi_{2}$ $=m_{2}l_{c2}g\cos\left(\theta_{1}+\theta_{2}-\frac{\pi}{2}\right)$ (17)
with parameters of the system as shown in Table 4. In the acrobot experiment that requires two control forces, the state equation of $\dot{\theta}_{1}$ is adapted to
$\ddot{\theta}_{1}=\frac{(u_{2}-d_{2}\ddot{\theta}_{2}-\phi_{1})}{d_{1}}.$
In practice, we convert this system to a stochastic first-order system
$d\mathbf{x}=f(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$ (18)
In every experiment, $\mathbf{R}\in\mathbb{R}^{C\times C}$ is set to $0.01\mathbf{I}$, $x^{\ast}$ to 1.5 and the initial states are sampled from $U(-0.1,0.1)$. Observations of $\theta_{i}$ are wrapped between [-$\pi$, $\pi$] and a trial terminates when $\dot{\theta_{1}}>8\pi$ or $\dot{\theta_{2}}>18\pi$. Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. In the third experiment, the policies can apply independent forces to both links, and the angular velocities are observed again. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=0.3\mathbf{I}$ to add observation noise. In experiment with partial state observability, $\mathbf{D}$ is changed to exclude the unobserved variables and $\mathbf{\Sigma}$ is set to the correct dimension.
@@ Appendix C Acrobot
Appendix D Continuous stirred tank reactor
The dynamics of the CSTR are defined as in [antonelli2003continuous], extended with a Wiener process to incorporate stochasticity. Let $\mathbf{x}=(c,T_{r},T_{c})^{\top}$. The CSTR is defined as a stochastic first-order system
$d\mathbf{x}=f(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$ (20)
with $\mathbf{V}=\textrm{diag}(0.025,3,3)$. The state equation is defined as
$f(\mathbf{x},\mathbf{u})=[f_{1}(\mathbf{x},\mathbf{u}),f_{2}(\mathbf{x},\mathbf{u}),f_{3}(\mathbf{x},\mathbf{u})]^{\top}$ (21)
with
$f_{1}(\mathbf{x},\mathbf{u})$ $=\frac{q_{r}}{V_{r}}\left(c_{f}-c\right)+k(T_{r})c$ (22)
$f_{2}(\mathbf{x},\mathbf{u})$ $=\frac{q_{r}}{V_{r}}\left(T_{f}-T_{r}\right)+\frac{-\Delta H}{\rho C_{p}}k(T_{r})c+\frac{UA}{\rho C_{p}V_{r}}\left(T_{c}-T_{r}\right)$ (23)
$f_{3}(\mathbf{x},\mathbf{u})$ $=\frac{q_{c}}{V_{c}}\left(T_{cf}-T_{c}\right)+\frac{UA}{\rho C_{p}V_{r}}\left(T_{r}-T_{c}\right).$ (24)
The descriptions of each parameter in the CSTR dynamics are shown in Table 5. Table 5 also contains the value, initial condition or sample range for each parameter. The variables that are indicated with a range are uniformly sampled within this range in every trajectory. The initial conditions of $c$, $T_{r}$ and $T_{c}$ are uniformly sampled from the given distribution. The control signal is clipped between 0 and 300. The reaction rate is defined as $k(T_{r})=k_{0}\exp\left(-{E}/{RT_{r}}\right)$.
where
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature $T_{r}^{\ast}$, while also minimising the control, i.e. the coolant flowrate.
Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table III). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid the policies in dealing with the values of the temperatures, which typically range between 200 and 600. To generate the observations, $\mathbf{D}$ is defined such that the concentration is excluded from the observations and $\mathbf{\Sigma}=7.5\mathbf{I}$.